If lim f(x) and lim g(x) exist as x approaches a then lim [ f(x) / g(x) ] = lim f(x) / lim g(x) as x approaches a. true or false


1
Expert's answer
2021-12-02T07:08:55-0500

False.

Correct statement

If lim "\lim f(x)" and "\lim g(x)" exist as "x" approaches "a" then


"\lim\limits_{x\to a}\dfrac{f(x)}{g(x)}=\dfrac{\lim\limits_{x\to a}f(x)}{\lim\limits_{x\to a}g(x)}\text{ if }\lim\limits_{x\to a}g(x)\not=0"